Hope someone can help me with this.
a. Find the vertex of the parabola.
f(x) = -x2 + 9x - 5
Which is the correct solution? Show work.

1. (-9/2; -263/4)
2. (9,-5)
3. (9/2,61/4)
4. (9/2,101/4)






b. Simplify (3 – 2i)(4 – 5i) and write the answer in the form a + bi, where a and b are real numbers. Show some work.







1. Solve the equation 5(4x – 1) – 2(x – 5) = 6(x + 2). Show some work.

Answer by scott8148(6628)

a. the vertex lies on the axis of symmetry __ the general equation is x=-b/(2a)

x=-9/[2(-1)] __ x=9/2 __ substituting __ f(9/2)=-(9/2)^2+9(9/2)-5 __ f(9/2)=-(81/4)+(81/2)-5 __ f(9/2)=61/4


b. FOILing __ 12-15i-8i+10i^2 __ 12-23i+10(-1) __ 2-23i


1. distributing __ 20x-5-2x+10=6x+12 __ 18x+5=6x+12 __ subtracting 6x+5 __ 12x=7 __ dividing by 12 __ x=7/12